# Open Agent Specification (Agent Spec) Technical Report

### v3

@@ 2.4 How Agent Spec Works Across Frameworks
Agent Spec serves as an abstraction layer above framework-specific implementations, acting as a unifying layer that encapsulates agent functionality beyond individual framework constraints.
Hence, the goal of Agent Spec is to define behavioral patterns for the different components that should be implemented according to the specificities of the different agentic frameworks. An Agent Spec runtime for a specific framework should be able to read the Agent Specification of a component (e.g., a Flow or an Agent), and then execute its logic, adhering to the behavior described in the Agent Spec.
As an alternative, if the agentic framework - used to implement the runtime - offers a serialization mechanism, it is also possible to build a translation script between the specification of an agent and the equivalent framework specific representation, then execute the runtime on the latter.
@@ 2.5 Implementation Details
2.5.1 Agent Spec SDKs
To facilitate the process of building programmatically framework-agnostic agents, Agent Spec SDK packages can be conveniently implemented. These SDKs should provide two functionalities:
  • Building Agent Spec component abstractions through the implementation of the relevant interfaces, in conformity with the Agent Spec.
  • Import and export these abstractions from and to their respective serialized version in JSON format.
As part of the Agent Spec effort, we provide Agent Spec’s Python SDK that we call PyAgentSpec. PyAgentSpec allows building Agent Spec-compliant agents in python. Users can create their own assistants using component classes that replicate the interface and behavior defined by Agent Spec, which can finally be exported in JSON format. Here is an example of how the definition of a simple agentic Flow prompting an LLM with a custom user input and returning the LLM’s output would look like in PyAgentSpec.
1from pyagentspec.property import Property
2from pyagentspec.flows.flow import Flow
3from pyagentspec.flows.edges import ControlFlowEdge, DataFlowEdge
4from pyagentspec.flows.nodes import LlmNode, StartNode, EndNode
5from pyagentspec.llms import VllmConfig
6llm_config = VllmConfig(
7 name="<Your Model Name Here>",
8 url="<url.of.your.llm.deployment:port>",
9 model_id="<provider/model-identifier>",
10)
11prompt_property = Property(
12 json_schema={"title": "prompt", "type": "string"}
13)
14llm_output_property = Property(
15 json_schema={"title": "llm_output", "type": "string"}
16)
17start_node = StartNode(name="start", inputs=[prompt_property])
18end_node = EndNode(name="end", outputs=[llm_output_property])
19llm_node = LlmNode(
20 name="simple llm node",
21 llm_config=llm_config,
22 prompt_template="{{prompt}}",
23 inputs=[prompt_property],
24 outputs=[llm_output_property],
25)
26flow = Flow(
27 name="Simple prompting flow",
28 start_node=start_node,
29 nodes=[start_node, llm_node, end_node],
30 control_flow_connections=[
31 ControlFlowEdge(name="start_to_llm", from_node=start_node, to_node=llm_node),
32 ControlFlowEdge(name="llm_to_end", from_node=llm_node, to_node=end_node),
33 ],
34 data_flow_connections=[
35 DataFlowEdge(
36 name="prompt_edge",
37 source_node=start_node,
38 source_output="prompt",
39 destination_node=llm_node,
40 destination_input="prompt",
41 ),
42 DataFlowEdge(
43 name="llm_output_edge",
44 source_node=llm_node,
45 source_output="llm_output",
46 destination_node=end_node,
47 destination_input="llm_output"
48 ),
49 ],
50)
To export the Flow that was just built in JSON format, the code is the following.
1serializer = AgentSpecSerializer()
2serialized_flow = serializer.to_json(flow)
The full documentation of PyAgentSpec, including installation instructions, APIs documentation and examples, is available at https://oracle.github.io/agent-spec/index.html.
2.5.2 Agent Spec SDK Plugins
The PyAgentSpec SDK also supports the concept of plugins, because we expect that teams using Agent Spec may want to create new component types to be included in their configurations. To this end, a plugin can enable newer components in addition to the standard components of Agent Spec, simply by specifying the list of component types that the plugin includes and by providing the serialization and deserialization logic for the plugin. Such plugins can be implemented for components not yet included in the first version of Agent Spec, such as multi-agent patterns, or planning and memory components. See below the abstract interface to be implemented for a serialization plugin (the deserialization interface follows the same logic).
1from abc import ABC, abstractmethod
2from typing import Any, Dict, List
3
4from pyagentspec.component import Component
5from pyagentspec.serialization.serializationcontext import SerializationContext
6
7class ComponentSerializationPlugin(ABC):
8 """Base class for Component serialization plugins."""
9
10 @property
11 @abstractmethod
12 def plugin_name(self) -> str:
13 """Return the plugin name."""
14 pass
15
16 @property
17 @abstractmethod
18 def plugin_version(self) -> str:
19 """Return the plugin version."""
20 pass
21
22 @abstractmethod
23 def supported_component_types(self) -> List[str]:
24 """Indicate what component types the plugin supports."""
25 pass
26
27 @abstractmethod
28 def serialize(
29 self, component: Component, serialization_context: SerializationContext
30 ) -> Dict[str, Any]:
31 """Serialize a component that the plugin should support."""
32 pass
2.5.3 Agent Spec Runtime Adapters
A runtime implementation makes Agent Spec components abstractions runnable by implementing the behavioral logic described in the Agent Spec language.
As previously discussed, there are several ways to implement runtimes for Agent Spec. Ideally, a runtime implementation is an agentic framework that natively supports Agent Spec: it should read an Agent Spec JSON specification and transform it directly into an equivalent component that can be executed.
As an example of an Agent Spec runtime implementation, we provide a package called WayFlow, which is a powerful, intuitive Python library for building sophisticated AI-powered assistants. It includes a library of plan steps to streamline the creation of AI-powered assistants, supports reusability, and it is ideal for rapid development. The WayFlow framework provides a serialization and deserialization interface to import and export Agent components according to the Agent Spec. It is a reference framework for Agent Spec and our runtime of reference, with native support for all Agent Spec Agents and Flows.
When an agentic framework does not allow reading Agent Specifications directly, a translation layer that transforms the Agent Specifications into a framework-specific representation could be adopted. This layer, which we can call Agent Spec Runtime adapter, loads an Agent Specification and transforms it into an equivalent component of the specific agentic framework.
Runtime adapters enable existing agentic frameworks that do not support Agent Spec natively executing Agent Specifications. For example, we also provide runtime adapters for some of the most popular agentic frameworks, including, but not limited to, LangGraph and Microsoft AutoGen. These adapters load Agent Specifications taking advantage of the Agent Spec Python SDK, and they transform the Agent Spec components into the respective versions defined in the agentic framework of interest.
2.6 Performance Evaluation & Benchmarks
Agent Spec enables key aspects of productizing agent-driven solutions, including comparison and evaluation of agent task performance across different executing frameworks and models, as well as optimizations (including costs) of execution for underlying hardware and frameworks. By covering commonly accepted concepts in agentic systems, such as workflows and agents, Agent Spec allows developers to map these concepts to various execution frameworks, facilitating comparison of framework specific task performance. Compute optimization is also possible, as Agent Spec enables hardware-optimized and framework-optimized implementation of agentic design patterns.
2.6.1 Portability Analysis
Agentic AI is rapidly evolving, and a unified benchmark for assessing agent reliability, particularly in enterprise settings, is currently lacking. Future updates for Agent Spec report will include empirical results demonstrating the reliability of Agent Spec-defined agents across different frameworks.
https://oracle.github.io/agent-spec/howtoguides/howto_execute_agentspec_across_frameworks.html provides an example of how an agent defined with Agent Spec can be run in different frameworks. In this Retrieval Augmented Generation case study, an agent utilizing the same tools is instantiated in AutoGen, LangGraph, and WayFlow. In a practical setting, a specific mapping from Agent Spec to each framework would be implemented.
2.6.2 Use Cases & Real-World Applications
A common use case of Agent Spec is the porting of agent-based assistants from one framework to another. Let’s take as an example a team that implemented agent-based workflows using AutoGen. As business and process needs to change, evaluations are conducted to consider alternative frameworks. Due to AutoGen’s specific architecture, migrating to another agentic framework proves to be complex with significant rewriting overhead. In this example, Agent Spec enables the definition of business unit-specific agentic systems using a declarative representation. This allows for conversion to specific framework formats or flexible execution across different agentic frameworks with appropriate adapters, enabling straightforward comparison of task performance and runtime cost depending on the execution framework. Consequently, utilizing Agent Spec throughout the agent design and deployment process reduces complexity, costs, implementation effort, and potential inconsistencies.
Another use case involves developing and deploying Agents across varied compute environments, such as cloud platforms and on-premise datacenters. In the cloud, Agents may run virtualized or serverless as part of a scalable system design. In an on-premise scenario, they may run within an application server. Agent Spec’s framework and runtime agnosticism enables development of the Agent in a local development environment, followed by flexible deployment across either of these scenarios.
A different compelling real-world application involves advanced multi-agent systems in the domain of financial crime compliance. By integrating generative AI agents, it enables to automate key investigative processes - supplementing human investigators with narrative-driven analyses, uncovering critical insights, and recommending evidence-based actions. In this context, a comprehensive Anti-Money Laundering platform may comprise specialized agents tasked with sanctions screening, transaction monitoring for suspicious activity, and the generation of enhanced due diligence reports. These sub-agents are often developed by independent teams using a variety of frameworks and technologies, and may need to be deployed across heterogeneous environments, including cloud, on-premises, or even within database systems. Leveraging Agent Spec’s framework-agnostic and declarative approach, development teams can seamlessly integrate and orchestrate these diverse agent workflows, reducing integration overhead and enabling consistent, scalable deployment. This holistic and automated strategy minimizes manual effort, decreases operational risk, and empowers financial institutions to respond more rapidly and consistently to evolving financial crime patterns, regardless of the underlying deployment environment.
2.7 Positioning in the Ecosystem
Agent Spec aims to streamline the architecture and design of agentic assistants and workflows, serving as a unified representation that abstracts away implementation details of specific agentic frameworks. It is a portable configuration language that describes agentic design patterns, components and expected behavior. However, Agent Spec is not the only effort aimed at unifying the different parts that compose a common agentic ecosystem.
  • The Model Context Protocol (MCP) [1], introduced by Anthropic, standardizes resource and data provision to agents. Resources are made available with a client/server based RESTful API.
  • Google’s Agent2Agent Protocol [6] and BeeAi’s Agent Communication Protocol (ACP) [3] propose standardized APIs for distributed agent communication.
  • Nvidia recently announced NeMo Agent toolkit which is a framework-independent library that treats “every agent, tool, and agentic workflow as a function call—enabling composability between these agents, tools, and workflows”[5]. The idea is to compose existing agentic solutions and resources, potentially written using different frameworks, into new systems.
While protocols like MCP and A2A standardize tool or resource provisioning as well as inter-agent communication, Agent Spec complements these efforts by enabling standardized configuration of components related to agentic system design and execution in general, as shown in Fig. 4. Hence, Agent Spec provides strong synergies with other standardization efforts through its abstract definition of agentic system design. In the context of ongoing unification efforts within the agentic system landscape, Agent Spec aims to serve as the "common foundation" that connects these initiatives, enhancing their effectiveness and fostering a more cohesive ecosystem. Support for the aforementioned protocols may be added in future versions of Agent Spec as new components.
Figure 4: Agent Spec complements other standardizations, such as MCP or A2A
2.8 Future Roadmap & Enhancements
2.8.1 Upcoming Features
Expanding the Open Agent Specification language We are looking forward to extending Agent Spec with new concepts and components in the upcoming versions of Agent Spec. A few notable ones include:
  • Memory
  • Planning
  • Datastores
  • Support for remote agents (e.g., A2A agents)
  • Support for new types of tools (e.g., MCP tools)
Extending runtimes support
We want to foster the implementation of Agent Spec runtime adapters for the popular agentic frameworks in the vibrant AI community. At the same time, we aim to offer a conformance test suite for runtimes to ensure consistent behavior in accordance with the Agent Specification.
Improving user experience
Even though Agent Spec SDKs simplify the creation and validation of Agent Specifications, creating agents by writing code might be complex and error-prone. We are working on a Drag&Drop UI that will allow users to build their own agents in a visual interface and then export the Agent Specification to be used in the framework of their choice via a runtime adapter.
2.8.2 Community & Ecosystem Growth
We strongly believe Agent Spec will contribute to streamlining agent development and open-sourcing its specification to encourage community contributions. This will foster a rich ecosystem of interchangeable and reusable designs and tools for agentic AI development. Integration of Agent Spec by other third-party open-source projects as an agent design language and specification will benefit the open-source community, which is currently evolving and lacks established common unified representation. In summary, the key aspects of impact on community growth are:
  • Design patterns that can be easily shared & reused to collaboratively build increasingly complex solutions.
  • Enable open exchange of building blocks and agentic patterns.
  • Encourage contributions by the open-source community.
To help guide the evolution of the Agent Specification in alignment with its core design principles, we will establish a steering committee. Similar to ONNX principles, this committee will work closely with the community to review proposed contributions, foster open discussion, and publish versioned updates.
@@ 3.1 Evaluation methodology
We evaluate Agent Spec on three popular benchmarks: SimpleQA Verified [4], $\tau^{2}$-Bench [2] and BIRD-SQL [14]. For each benchmark, we select one or more Agent Spec JSON configurations and execute the same specification across four framework runtimes: AutoGen, CrewAI, LangGraph, and WayFlow. This approach allows direct comparison of behavioral consistency, correctness, and latency across heterogeneous frameworks. In this evaluation, the same agent utilizing identical tools is instantiated in AutoGen, CrewAI, LangGraph, and WayFlow.
@@
We evaluate on the full set of 1,000 questions using two representative assistants: (i) a ReAct-style agent equipped with a web-browsing tool, and (ii) a flow-based agentic RAG solution that applies question decomposition, web-browsing, and self-reflection to answer complex user queries [^1] Note that the browse tool makes the task [4] easier, but it is acceptable since we are demonstrating Agent Spec’s functional effectiveness instead of competing on the benchmark leaderboard.. We also include the performance of a single LLM call as a baseline.
@@
We evaluate execution accuracy on the 1,534 questions from BIRD-SQL dev with two types of assistants. The first assistant is a ReAct-style agent equipped with a tool to validate that a query can be executed, and a "thinking" tool to give opportunities to the agent to think between tool calls. The second assistant is a flow-based assistant composed of a planning step, a step to generate the SQL query and a final post-processing step to self-reflect on the generated query. We also include the performance of a single LLM call as a baseline.
@@ 3.2 Results and Insights
4 Contributors and Acknowledgements
Please cite this work as “Open Agent Specification (2025)”.
Yassine Benajiba, Cesare Bernardis, Vladislav Blinov, Paul Cayet, Hassan Chafi, Abderrahim Fathan, Louis Faucon, Damien Hilloulin, Sungpack Hong, Ingo Kossyk, Rhicheek Patra, Sujith Ravi, Jonas Schweizer, Jyotika Singh, Shailender Singh, Xuelin Situ, Weiyi Sun, Kartik Talamadupula, Jerry Xu, Ying Xu. [^2] Author list is sorted alphabetically.
We are also thankful to our Oracle colleagues who engaged in active discussions during the design of Agent Spec and provided valuable feedback that helped us refine its specifications.

### v4

@@ 4.4 How Agent Spec Works Across Frameworks
The primary objective of Agent Spec is to define behavioral patterns for the various components that can be implemented according to the specific characteristics of different agentic frameworks. A runtime environment developed for a particular framework should be capable of reading the serialized configuration of a component (such as a Flow or an Agent) and executing its logic in accordance with the behaviors prescribed by Agent Spec. Every Agent Spec configuration includes a top-level agentspec_version property specifying the specification version it follows.
Alternatively, if the agentic framework used to implement the runtime provides its own serialization mechanism, a translation layer could be introduced. This layer can convert an Agent Spec configuration into the equivalent framework-specific representation, allowing the runtime to execute the agent directly using the translated format.
@@ 4.5 Implementation Details
To facilitate the programmatic construction of framework-agnostic agents, Agent Spec supports the development of dedicated Software Development Kits (SDKs). Our reference SDK, PyAgentSpec, also includes a plugin architecture that allows teams adopting Agent Spec to define and integrate new component types within their configurations as needed. A runtime adapter implementation makes Agent Spec component abstractions executable by providing the behavioral logic defined in the Agent Spec language. Beyond simple translation, these runtime adapters act as “compiler” for Agent Spec, converting its declarative specifications into executables for target frameworks and bridges high-level declarative design with low-level execution semantics. Please see Appendix A for an example of an agentic flow in PyAgentSpec and Appendix B for details on the reference Agent Spec SDK, PyAgentSpec, the plugin architecture and runtime adapters.
@@ 5.1 Evaluation methodology
We evaluate Agent Spec on three popular benchmarks: SimpleQA Verified [haas2025simpleqaverified], $\tau^{2}$-Bench [barres2025tau2] and BIRD-SQL [li2024can]. For each benchmark, we select one or more Agent Spec JSON configurations and execute the same specification across four framework runtimes: AutoGen, CrewAI, LangGraph, and WayFlow. This approach allows direct comparison of behavioral consistency, correctness, and latency across heterogeneous frameworks. In this evaluation, the same agent utilizing identical tools is instantiated in AutoGen, CrewAI, LangGraph, and WayFlow.
@@
We evaluate on the full set of 1,000 questions using two representative assistants: (i) a ReAct-style agent equipped with a web-browsing tool, and (ii) a flow-based agentic RAG solution  [^3] Details about the flow is provided in Appendix D.1. that applies question decomposition, web-browsing, and self-reflection to answer complex user queries [^4] Note that the browse tool makes the task [haas2025simpleqaverified] easier, but it is acceptable since we are demonstrating Agent Spec’s functional effectiveness instead of competing on the benchmark leaderboard.. We also include the performance of a single LLM call as a baseline.
@@
We evaluate execution accuracy on the 1,534 questions from BIRD-SQL dev with two types of assistants. The first assistant is a ReAct-style agent equipped with a tool to validate that a query can be executed, and a "thinking" tool to give opportunities to the agent to think between tool calls. The second assistant is a flow-based assistant  [^5] Details about the flow is provided in Appendix D.2. composed of a planning step, a step to generate the SQL query and a final post-processing step to self-reflect on the generated query. We also include the performance of a single LLM call as a baseline.
@@ 5.2 Results and Insights
6 Conclusion & Potential Enhancements
This paper introduced Agent Spec, a declarative, framework-agnostic specification for defining AI agents and workflows. By decoupling design from execution, it enables define-once, run-anywhere portability and interoperability across agentic frameworks. We formalized its core components and semantics, and released toolsets: a Python SDK (PyAgentSpec), a reference runtime (WayFlow), and adapters for frameworks (e.g., LangGraph and AutoGen).
We empirically show that Agent Spec serves as an Evaluation harness, enabling consistent comparison of agent and flow patterns across frameworks. We outline potential extensions to Agent Spec’s reusability, interoperability, and developer experience in Appendix E. Planned enhancements include compiler-style optimizations such as fusing consecutive LLMNodes and using structured generation to reduce redundant calls. Looking ahead, we envision Agent Spec as a foundation for standardizing agent design & evaluation.
Appendix A Simple Agentic Flow in PyAgentSpec
The following example illustrates the definition of a simple agentic Flow in PyAgentSpec that prompts an LLM with a user-provided input and returns the model’s generated output.
from pyagentspec.property import Property
from pyagentspec.flows.flow import Flow
from pyagentspec.flows.edges import ControlFlowEdge, DataFlowEdge
from pyagentspec.flows.nodes import LlmNode, StartNode, EndNode
from pyagentspec.llms import VllmConfig
llm_config = VllmConfig(
name="<Your Model Name Here>",
url="<url.of.your.llm.deployment:port>",
model_id="<provider/model-identifier>",
prompt_property = Property(
json_schema={"title": "prompt", "type": "string"}
llm_output_property = Property(
json_schema={"title": "llm_output", "type": "string"}
start_node = StartNode(name="start", inputs=[prompt_property])
end_node = EndNode(name="end", outputs=[llm_output_property])
llm_node = LlmNode(
name="simple llm node",
llm_config=llm_config,
prompt_template="{{prompt}}",
inputs=[prompt_property],
outputs=[llm_output_property],
flow = Flow(
name="Simple prompting flow",
start_node=start_node,
nodes=[start_node, llm_node, end_node],
control_flow_connections=[
ControlFlowEdge(name="start_to_llm", from_node=start_node, to_node=llm_node),
ControlFlowEdge(name="llm_to_end", from_node=llm_node, to_node=end_node),
data_flow_connections=[
DataFlowEdge(
name="prompt_edge",
source_node=start_node,
source_output="prompt",
destination_node=llm_node,
destination_input="prompt",
DataFlowEdge(
name="llm_output_edge",
source_node=llm_node,
source_output="llm_output",
destination_node=end_node,
destination_input="llm_output"
Listing 1: Simple Agentic Flow in PyAgentSpec
To export the Flow that was just created in JSON format, use the following code.
serializer = AgentSpecSerializer()
serialized_flow = serializer.to_json(flow)
Listing 2: Export flow to JSON
Appendix B Appendix - Implementation Details
B.1 Agent Spec SDKs
Agent Spec SDKs are expected to provide two core functionalities:
  • Construction of Agent Spec component abstractions through the implementation of the corresponding interfaces, in full conformity with the Agent Spec standard.
  • Import and export of these abstractions to and from their serialized representations in JSON format.
As part of the Agent Spec initiative, we provide a reference implementation in Python, called PyAgentSpec [pyagentspec]. PyAgentSpec enables developers to construct Agent Spec-compliant agents directly in Python. Users can define their own assistants by instantiating component classes that mirror the interfaces and behaviors prescribed by Agent Spec, and subsequently export these definitions to JSON.
B.2 Agent Spec SDK Plugins
A plugin can extend the standard set of Agent Spec components by defining additional component types and specifying the corresponding serialization and deserialization logic required to handle them.
These plugins enable support for components not yet included in the current version of Agent Spec, such as those related to memory.
The abstract interface for implementing a serialization plugin is shown below; the deserialization interface follows the same design principles.
from abc import ABC, abstractmethod
from typing import Any, Dict, List
from pyagentspec.component import Component
from pyagentspec.serialization.serializationcontext import SerializationContext
class ComponentSerializationPlugin(ABC):
"""Base class for Component serialization plugins."""
@property
@abstractmethod
def plugin_name(self) -> str:
"""Return the plugin name."""
pass
@property
@abstractmethod
def plugin_version(self) -> str:
"""Return the plugin version."""
pass
@abstractmethod
def supported_component_types(self) -> List[str]:
"""Indicate what component types the plugin supports."""
pass
@abstractmethod
def serialize(
self, component: Component, serialization_context: SerializationContext
) -> Dict[str, Any]:
"""Serialize a component that the plugin should support."""
pass
Listing 3: Interface for serializing an Agent Spec plugin
B.3 Agent Spec Runtime Adapters
There are multiple ways to implement runtimes for Agent Spec. Ideally, a runtime corresponds to an agentic framework that natively supports Agent Spec, i.e., it can directly read an Agent Spec JSON specification and instantiate an equivalent executable component.
As a reference example, we provide a runtime implementation called WayFlow [^6] https://github.com/oracle/wayflow, a powerful and intuitive Python library for constructing sophisticated AI-powered assistants. WayFlow includes a library of prebuilt plan steps to streamline assistant creation, promotes component reusability, and supports rapid prototyping. The framework implements both serialization and deserialization interfaces, enabling the import and export of Agent components that conform to Agent Spec. It serves as the reference runtime for Agent Spec, offering native support for all Agent Spec Agents and Flows.
In cases where an agentic framework does not natively support Agent Spec, a translation layer that transforms the Agent Specifications into a framework-specific representation could be adopted. This layer, which we call Agent Spec Runtime Adapter, loads an Agent Specification and converts it into an equivalent representation compatible with the target agentic framework, that can be finally executed.
We provide runtime adapters [^7] https://github.com/oracle/agent-spec/tree/main/adapters for some of the most popular agentic frameworks, including, but not limited to, LangGraph and Microsoft AutoGen. These adapters leverage the Agent Spec Python SDK to load specifications and translate Agent Spec components into their respective framework-specific equivalents.
Appendix C Appendix - Use Cases & Applications
A common use case for Agent Spec is the portability of agent-based assistants across different frameworks. Consider, for instance, a team that has implemented agent-based workflows using AutoGen. As business requirements and process needs evolve (or framework is deprecated [autogenmigration]), the team evaluates alternative frameworks. However, due to AutoGen’s specific architecture, migrating these workflows to another agentic framework can be complex and typically entails substantial rewriting overhead.
In such cases, Agent Spec enables the development of agentic systems, tailored to specific business units, using a declarative representation. This representation can then be converted into framework-specific formats or executed flexibly across multiple frameworks via appropriate runtime adapters. This approach facilitates direct comparison of task performance and execution costs across frameworks. Consequently, employing Agent Spec throughout the agent design and deployment lifecycle significantly reduces complexity, implementation effort, maintenance cost, and potential inconsistencies.
Another key use case involves the development and deployment of Agents across heterogeneous compute environments, such as cloud platforms and on-premises data centers. In cloud-based systems, Agents might run in virtualized or serverless configurations as part of a scalable architecture, whereas in on-premises contexts, they might operate within traditional application servers. Agent Spec’s framework and runtime-agnostic design enables development of the Agent in a local development environment, followed by flexible deployment across either of these scenarios.
An illustrative example of this interoperability is available at https://oracle.github.io/agent-spec/howtoguides/howto_execute_agentspec_across_frameworks.html which demonstrates how an agent defined with Agent Spec can be executed across different frameworks.
Appendix D Appendix - Agentic Flows
D.1 RAG Agentic Flow
The Flow-based Agentic RAG used for SimpleQA is an advanced question-answering system that adapts its reasoning strategy to the complexity of each query. The assistant begins with a basic Retrieval-Augmented Generation (RAG) pipeline for straightforward questions and routes to iterative, modular reasoning involving question decomposition, search, result deduplication, and answer synthesis for more complex or open-ended queries.
Initially, the user’s query is embedded and sent to a web search tool, which retrieves a set of relevant results. These results are synthesized into a preliminary answer, which is judged by an LLM to assess if it is likely sufficient to answer the initial user question. If the system determines that this initial response adequately answers the user’s question, it is returned immediately, minimizing latency.
If the response is missing, incomplete or lacks confidence, the system automatically transitions to a more sophisticated, iterative research mode. This pipeline begins by constructing a comprehensive context from the user’s original question, any prior sub-questions, retrieved information, and feedback from earlier cycles. It then decomposes the complex query into focused sub-questions, which helps target web search and filter out irrelevant details. For each sub-question, the system retrieves new results, deduplicates overlapping content, and reassesses previously gathered evidence. The assistant then synthesizes a detailed answer and an evaluation component determines whether the answer is now comprehensive; if so, it finalizes and returns the response. Otherwise, the refinement loop continues until the assistant achieves a confident response or a maximum number of iterations is reached.
D.2 NL2SQL Agentic Flow
This flow incorporates a structured sequence of agentic components to improve reasoning and reliability. It begins with an LLM planning node that interprets the user’s query and formulates a high-level strategy. The plan is then executed by a ReAct-style agent equipped with external tools for retrieval, computation, or other actions. Finally, a post-processing LLM node reviews and refines the agent’s output to verify correctness, check for specified errors, and format the final response before returning it to the user. This design was introduced because vanilla ReAct agents based on GPT-4.1 often fail to autonomously generate explicit plans or to perform final verification steps, leading to inconsistent or incomplete results.
Appendix E Appendix - Potential Enhancements
E.1 Areas of Potential Extension
Several extensions could broaden Agent Spec’s applicability to common agentic patterns and enterprise-scale requirements while maintaining its declarative and framework-agnostic design principles. Key areas include:
  • Memory components: Representations for short- and long-term memory, including context persistence and retrieval strategies.
  • Planning modules: Declarative definitions of planner components for hierarchical or goal-oriented task decomposition.
  • Datastore interfaces: Native integration with datastores.
  • Remote agents: Support for defining and communicating with remote agents (e.g., using A2A).
Each of these enhancements would continue to adhere to the same design principles of clear semantics, explicit I/O, and declarative configuration, while staying framework-agnostic.
E.2 Runtime Coverage and Conformance
Expanding runtime coverage remains a core objective. We encourage community-led development of runtime adapters for additional agentic frameworks, along with contributions to an extended conformance test suite. A more robust test suite will help ensure that identical Agent Spec configurations yield consistent behaviors across diverse runtime environments, thereby improving reproducibility, reliability, and interoperability.
E.3 Improving User Experience
While the existing SDKs simplify authoring and validating Agent Specifications, future efforts will focus on improving usability and accessibility, particularly for non-programmatic users. Potential directions include:
  • Visual authoring tools: Drag-and-drop interfaces to construct agents and export Agent Spec configurations to be used in the preferred framework via a runtime adapter.
  • Advanced diagnostics: Enhanced schema validation, error explanations, and debugging utilities within SDKs.
These efforts aim to make Agent Spec more accessible for practitioners, educators, and enterprise developers alike.